A beam of Monochromatic light of wavelength 582 nm falls normally on a glass wedge with the wedge angle 20'' of arc. If the refractive index of the glass is 1.5, find the number of dark fringes per cm of the wedge length.
1
Expert's answer
2019-04-02T10:55:54-0400

The condition for interference is


"{2}{t}={m}{\lambda} (1)"

where t is the thickness of the air wedge at a distance x from the point of contact of the glass plates.


θ is the angle between the plates in radians (this angle is small, so tan θ = θ in radians). In this case, we got



"\theta=\frac{t}{x} (2)"

So


"2{\theta}{x}={m}{\lambda} (3)"

We got



"\frac{m}{x}=\frac{2\theta}{\lambda} (4)"

20'' is equal to 9.59×10-5 radians


Using (4) we get 3.29 fringes/cm


Answer:

3.29 fringes/cm